OpenStudy (blake57roger):

Find the simple interest. principal = $750; rate = 14%; time = 4.5 years

OpenStudy (gorv):

\[simple interest = P*R*T\]

OpenStudy (gorv):

p=750 r=14/100 t=4.5

OpenStudy (blake57roger):

A. $105 B. $420 C. $472.50 D. $1222.50

OpenStudy (blake57roger):

is it c

OpenStudy (whpalmer4):

what is \[750*\frac{14}{100}*4.5=\]?

OpenStudy (blake57roger):

Based on simple interest, find the total amount saved or owed. principal = $1056; rate = 3.75%; time = 6 years

OpenStudy (whpalmer4):

the answer is C if (and only if) that number is 472.50.

OpenStudy (whpalmer4):

Okay, you've seen how this works. P = 1056, r = 3.75/100, t = 6 What is the answer?